# Towards Black-Box Membership Inference Attack for Diffusion Models

### v1

@@ Abstract
Identifying whether an artwork was used to train a diffusion model is an important research topic, given the rising popularity of AI-generated art and the associated copyright concerns. The work approaches this problem from the membership inference attack (MIA) perspective. We first identify the limitations of applying existing MIA methods for copyright protection: the required access of internal U-nets and the choice of non-member datasets for evaluation. To address the above problems, we introduce a novel black-box membership inference attack method that operates without needing access to the model’s internal U-net. We then construct a DALL-E generated dataset for a more comprehensive evaluation. We validate our method across various setups, and our experimental results outperform previous works.
@@ 1 Introduction
Recently, there has been a surge in the popularity of generative models, with diffusion models in particular, gaining huge attention within the AI community [38, 41, 42]. These models have demonstrated remarkable capabilities across various tasks, including unconditional image generation [12, 40], text-to-image generation [32, 49, 29] and image-to-image generation [33]. This surge has given rise to powerful AI art models such as DALL-E 2 [31], Stable Diffusion [32] and Imagen [34]. The field of AI-generated art holds a promising future and is expected to have widespread impact.
Effective training of diffusion models requires high-quality data. It is thus crucial to design an algorithm that can identify whether a specific artwork has been used during the training of a model, thereby providing protection for these artworks and detecting misuse of data. This is a classic problem in the field of machine learning, first introduced by [37] and named “membership inference attack”.
@@
Despite great progress, previous methods are not yet ready for copyright protection in diffusion models for two reasons. First, most existing approaches heavily rely on checking whether the U-net of the model predicts noise accurately, which is not practical since most commercial diffusion models available today offer only API access, while the U-net remains hidden. We refer to MIA with access only to generated images as the Black-Box MIA. Second, existing evaluations for stable diffusion models primarily focus on content shifts in the data, while the shifts in art styles are not covered (illustrated in Figure 3).
To address the above issues, we propose a black-box membership inference attack method that does not require access to the U-net of the model. We posit that if we alter an image using the target diffusion model’s variation API, the sampling process will be captured by a region of “attraction” if the model has seen this image during training (illustrated in Figure 2). By quantifying the disparity between the resulting reconstructed image with the original, we can effectively determine whether the image was part of the training dataset.
We also construct a new evaluation dataset named LAION by DALL-E reflecting style shifts in data rather than content shifts. In addition to evaluating our methods on existing datasets across multiple models, we test MIA on the stable diffusion model to classify LAION by DALL-E from the ground truth training data LAION-5B.
The main contributions of this work are listed as follows:
@@
    We propose a black-box membership inference attack method that does not require access to the model’s internal structure. Instead, we leverage the unique properties of diffusion models. Our method only involves using the model’s variation API to alter an image and compare it with the original one. We name our method ReDiffuse.
@@
    We evaluate our method using DDIM and Stable Diffusion models on classical datasets, including CIFAR10, STL10-Unlabeled [2], LAION-5B [36], etc. Our method is shown to be effective and robust across datasets, and it outperforms the previous methods.
@@
    We propose a new MIA evaluation dataset which is complementary to previous work. We use the image titles from LAION-5B as prompts for DALL-E’s API [31] to generate images of the same contents but different styles. Under this setup, previous membership inference attack algorithms are no longer effective, while our method remains robust.
@@ 2 Related Works
In the era of big data, preserving data privacy is paramount, especially in training diffusion models using sensitive datasets like artists’ artworks, which are protected by copyright laws. Membership inference attacks, initially introduced by [37], serve as an effective means to detect potential misuse of data without proper authorization. Its objective is to ascertain whether a particular data sample participated in the training phase of a target model. This approach is instrumental in probing privacy breaches and identifying illicit data utilization. Researchers primarily focus on membership inference attacks for classification models [35, 48, 24, 22], embedding models [39, 6, 26], and generative models [8, 10, 1].
In the domain of membership inference attacks against diffusion models, [47, 14] use a white-box approach, which assumes access to the entire diffusion model and utilizes loss and likelihood to determine whether a sample is in the training set. More recently, [5, 18, 43] have relaxed these requirements, eliminating the need for the entire model. They leverage the insight that samples within the training set yield more accurate noise predictions, thereby achieving high accuracy in membership inference attack tasks. However, they also require the outputs of the U-net, as it is necessary to obtain the noise predictions of intermediate steps.
@@ 3 Preliminary
In this section, we begin by discussing several popular diffusion models. We first introduce the Denoising Diffusion Probabilistic Model (DDPM) [12]. Then, we extend to the Denoising Diffusion Implicit Model (DDIM) [40], which is a variant of DDPM used to accelerate image generation. Lastly, we discuss the Stable Diffusion [32], a model that condenses images into a latent space and generates images grounded in text descriptions.
@@
$q(x_{t}\mid x_{t-1})$ $=\mathcal{N}(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}\mathbf{I})\,,$
$p_{\theta}(x_{t-1}\mid x_{t})$ $=\mathcal{N}(x_{t-1};\mu_{\theta}(x_{t},t),\Sigma_{\theta}(x_{t},t))\,,$
@@ 4 Algorithm Design
In this section, we introduce our algorithm. We begin by discussing the definition of black-box API and the limitations of previous membership inference attack methods. In our formulations we assume DDIM as our target model. The formulations for DDPM is highly similar and we omit it for brevity. We will discuss the generalization to the latent diffusion model in Section 5.
4.1 Towards Black-Box MIA for Diffusion Models
Most previous works on membership inference attacks against diffusion models aim to prevent data leakage, and hence can rely on thresholding the model’s training loss. For instance, [14] involves a direct comparison of image losses, while [5, 18] evaluates the accuracy of the model’s noise prediction at initial or intermediate steps. However, the required access to the model’s internal U-net structure prevent applications to copyright protection because most servers typically provide only black-box API access.
We denote a model API as black-box if it only requires the input image and returns the output image. Specifically, we do not have access to the internal structure of the model. Below we formulate the definition of the black-box variation API used in our algorithm.
Definition 1 (The black-box variation API).
@@ 4.1 Towards Black-Box MIA for Diffusion Models
This definition aligns with the image-to-image generation method of the diffusion model, making access to this API practical in many setups [25, 33, 45]. Some APIs provide the user with a choice of $t$, while others do not and use the default parameter. We will discuss the influence of different diffusion steps in Section 6.4, showing that the attack performances are relatively stable and not sensitive to the selection of $t$. We also note that the for the target model, we can substitute $\phi_{\theta}(x_{t},0)$ with other sampling methods, such as the Euler-Maruyama Method [27] or Variational Diffusion Models [16].
Figure 1: The overview of ReDiffuse. We independently input the image to the black-box variation API $n$ times with diffusion step $t$. We take the average of the output images and compare them with the original ones. If the difference is below a certain threshold, we determine that the image is in the training set.
@@ 4.2 Algorithm
In this section, we present our algorithm for black-box MIA. Our key intuition comes from the reverse SDE dynamics in continuous diffusion models [41, 42]. Specifically, one simplified form of the reverse SDE (i.e., the denoise step) has the formulation below:
$\dot{X_{t}}=(X_{t}/2-\nabla_{x}\log p(X_{t}))+dW_{t}\,,\quad t\in[0,T]\,.$ (3)
where $W_{t}$ denotes a standard Brownian motion (with abuse of notation $t$) and $T$ is the generation time.
Figure 2: The intuition of our algorithm design. We denote $x$ as the target image, $x_{i}$ as the $i$-th image generated by the variation API, and $\hat{x}$ as the average image of them. For member image $x$, the difference $v=x-\hat{x}$ will be smaller after averaging due to $x_{i}$ being an unbiased estimator.
For a training data $x$, a score-based model $s_{\theta}(x)$ can be trained to fit $\nabla_{x}\log p(x)$ for image generation. The distribution $x_{0}\sim p_{0}\approx\delta_{x}$ is a delta distribution supported on the trained image $x$.
Then for a given image $x$, we generate $n$ noisy images $\{y_{i}\}_{i\leq n}$, $y_{i}=\beta_{T}x+(1-\beta_{T})\epsilon_{i}$, and then denoise to get $n$ reconstructed image $\{\hat{x}_{i}\}_{i\leq n}$. If the score function $s_{\theta}(x)$ is learned properly, we will generate the image satisfying the following distribution:
$\hat{x}_{i}\sim\mathcal{N}(x,\sigma_{T}I),$
where $\sigma_{T}$ is a constant determined by the injected noise level $\beta_{T}$.
The key guarantee is that when the score function is learned for a data point $x$, then the reconstructed image $\hat{x}_{i}$ is an unbiased estimator of $x$. Hence, averaging over multiple independent samples $\hat{x}_{i}$ would greatly reduce the estimation error (see Theorem 1). On the other hand, for a non-member image $x^{\prime}$, the unbiasedness of the denoised image is not guaranteed. We illustrate the intuition in Figure 2.
With the above intuition, we introduce the details of our algorithm. We independently apply the black-box variation API $n$ times with our target image $x$ as input, average the output images, and then compare the average result $\hat{x}$ with the original image. We will discuss the impact of the averaging number $n$ in Section 6.4. We then evaluate the difference between the images using an indicator function:
$f(x)=\mathbf{1}[D(x,\hat{x})<\tau]\,.$
@@
Algorithm 1 Black-Box MIA through ReDiffuse
@@
4.3 Theoretical Analysis
[…]
@@ 4.3 Theoretical Analysis
Suppose the DDIM model can learn a parameter $\theta$ such that, for any $x\sim D_{\text{training}}$, the prediction error $\|\epsilon-\epsilon_{\theta}(\sqrt{\bar{\alpha}_{t}}x+\sqrt{1-\bar{\alpha}_{t}}\epsilon,t)\|$ is a centered random variable $X$ with finite cumulant-generating function [7]. We also assume the sampling interval $k$ is equal to the variation API diffusion step $t$. Let $\hat{x}$ be the average of $n$ output images of $x$ using the variation API. Then we have
$\mathbb{P}(\|\hat{x}-x\|\geq\beta)\leq\exp(-n\Psi^{*}_{X}(\frac{\beta\sqrt{\bar{\alpha}_{t}}}{\sqrt{1-\bar{\alpha}_{t}}}))\,,$
where $\beta>0$ and $\Psi^{*}_{X}$ is the Fenchel-Legendre dual of the cumulant-generating function $\Psi_{X}$.
@@
5 MIA on Latent Diffusion Models
In this section, we discuss how we can generalize our algorithm to latent diffusion models. We will focus on stable diffusion [32], as it is one of the most popular choices.
5.1 The Variation API and Modification to ReDiffuse
We note that the variation API for stable diffusion is different from DDIM, as it includes the encoder-decoder process. Again we denote $\hat{x}=V_{\theta}(x,t)$, and the details are as follows:
@@ 5.1 The Variation API and Modification to ReDiffuse
Then we introduce the modification of our algorithm in this setup. For DDPM or DDIM, we evaluate the difference between the reconstructed image and the original image. However, for images processed by the stable diffusion model, which are typically larger in size, directly comparing the restored image with the original may lead to drastic differences. To enhance our algorithm, we modify the decision function, by independently adding random noise to the original image twice and then comparing the differences between the two restored images $\hat{x}_{1}$ and $\hat{x}_{2}$:
$f(x)=\mathbf{1}[D(\hat{x}_{1},\hat{x}_{2})<\tau]\,.$
This modification has shown promising experimental results, as detailed in Section 6.3. We refer to this improved version as “ReDiffuse+”.
5.2 LAION by DALL-E: A Style-shift Evaluation Dataset
Figure 3: Comparison of data used in our new experiment versus the previous work. It shows that our dataset, LAION by DALL-E, achieves minimal content differences from the original LAION-5B images. In contrast, the COCO dataset images used in previous works show significant content differences from the LAION-5B images, which could potentially impact the evaluation results.
The dataset used for evaluation of stable diffusion in prior studies [5, 18] possesses certain limitations. In the context of attacking stable diffusion, LAION-5B served as the in-training-set data, while the COCO dataset was designated as the out-of-training-set data. However, the inherent content disparities between these two datasets are considerable, shown in Figure 3. Therefore, the MIA algorithm’s capability to differentiate between them may not solely stem from whether the data was employed during training, but also from their content disparities.
Complementary to previous work, we propose a new evaluation dataset aimed at making the out-of-training-set data contain the same content as the training data but with different styles. Concretely, we use the textual descriptions from the LAION-5B dataset (which was used for training) to generate new images with the DALL-E model. We refer to this new dataset as “LAION by DALL-E” (shown in Figure 3). These generated images will serve as images that the stable diffusion model has not seen during training. By including this dataset in the evaluation, we achieve a more comprehensive evaluation of MIA methods for stable diffusion.

### v2

@@ Abstract
Given the rising popularity of AI-generated art and the associated copyright concerns, identifying whether an artwork was used to train a diffusion model is an important research topic. The work approaches this problem from the membership inference attack (MIA) perspective. We first identify the limitation of applying existing MIA methods for proprietary diffusion models: the required access of internal U-nets. To address the above problem, we introduce a novel membership inference attack method that uses only the image-to-image variation API and operates without access to the model’s internal U-net. Our method is based on the intuition that the model can more easily obtain an unbiased noise prediction estimate for images from the training set. By applying the API multiple times to the target image, averaging the outputs, and comparing the result to the original image, our approach can classify whether a sample was part of the training set. We validate our method using DDIM and Stable Diffusion setups and further extend both our approach and existing algorithms to the Diffusion Transformer architecture. Our experimental results consistently outperform previous methods.
@@ 1 Introduction
Recently, there has been a surge in the popularity of generative models, with diffusion models in particular, gaining huge attention within the AI community [41, 44, 45]. These models have demonstrated remarkable capabilities across various tasks, including unconditional image generation [13, 43], text-to-image generation [35, 52, 30] and image-to-image generation [36]. This surge has given rise to powerful AI art models such as DALL-E 2 [34], Stable Diffusion [35], and Imagen [37]. AI-generated art holds a promising future and is expected to have a widespread impact.
Effective training of diffusion models requires high-quality data. It is thus crucial to design an algorithm that can identify whether a specific artwork has been used during the training of a model, thereby providing protection for these artworks and detecting misuse of data. This is especially important due to the rapid growth of generative models, which has raised concerns over intellectual property (IP) rights, data privacy, and the ethical implications of training on copyrighted or proprietary content without consent. As these models are increasingly deployed across industries, detecting whether a specific piece of content was used in training can help prevent unauthorized use of artistic works, protecting creators’ copyrights and ownership rights. This is a classic problem in the field of machine learning, first introduced by [40] and named “membership inference attack”.
@@
Despite great progress, previous methods are not yet ready for MIA in proprietary diffusion models. Most existing approaches heavily rely on checking whether the U-net of the model predicts noise accurately, which is not practical since most commercial diffusion models available today offer only API access, while the U-net remains hidden.
To address the above issue, we propose a membership inference attack method that relies on the variation API and does not require access to the denoise model (e.g., U-net). We observe that if we alter an image using the target diffusion model’s variation API, the sampling process will be captured by “a region of attraction” if the model has seen this image during training (illustrated in Figure 2). Based on the above observation, we propose the ReDiffuse algorithm for MIA with image-to-image variation API, and can detect member images without accessing the denoise model. Our main contributions are listed as follows:
@@
    We propose a membership inference attack method that does not require access to the model’s internal structure. Our method only involves using the model’s variation API to alter an image and compare it with the original one. We name our method ReDiffuse.
@@
    We evaluate our method using DDIM [43] and Stable Diffusion [35] models on classical datasets, including CIFAR10/100 [21], STL10-Unlabeled [2], LAION-5B [39], etc. Our method outperforms the previous methods.
@@
    We extend both existing algorithms and our own algorithm to the Diffusion Transformer [32] architecture, implementing the membership inference attack within this model framework for the first time. Experimental results demonstrate that our algorithm is consistently effective.
@@ 2 Related Works
In the era of big data, preserving data privacy is paramount. The training of diffusion models may involve sensitive datasets like artists’ artworks, which are protected by copyright laws. Membership inference attacks, initially introduced by [40], serve as an effective means to detect potential misuse of data without proper authorization. Its objective is to ascertain whether a particular data sample participated in the training phase of a target model. This approach is instrumental in probing privacy breaches and identifying illicit data utilization. Researchers primarily focus on membership inference attacks for classification models [38, 51, 25, 23], embedding models [42, 7, 27], and generative models [9, 11, 1].
In the domain of membership inference attacks against diffusion models, [50, 15] use a white-box approach, which assumes access to the entire diffusion model and utilizes loss and likelihood to determine whether a sample is in the training set. [6, 19, 46] have relaxed these requirements, eliminating the need for the entire model. They leverage the insight that samples within the training set yield more accurate noise predictions, thereby achieving high accuracy in membership inference attack tasks. However, they also require the outputs of the U-net, as it is necessary to obtain the noise predictions of intermediate steps. Recently, [31] proposed a black-box membership inference attack method against diffusion models. Their method identifies whether a specific image is in a finetuning dataset of 100 images by calculating the difference between the generated image and the target image, using the corresponding prompt as input. Their approach leverages the model’s tendency to memorize finetuning images and generate similar outputs. In contrast, we focus on detecting whether an image is in the pretraining dataset, where a pre-trained model often produces diverse outputs for the same prompt, making detection more challenging.
@@ 3 Preliminary
In this section, we begin by introducing the notations used for several popular diffusion models. We first introduce the Denoising Diffusion Probabilistic Model (DDPM) [13]. Then, we extend to the Denoising Diffusion Implicit Model (DDIM) [43] and Stable Diffusion [35], which are variants of DDPM used to accelerate image generation or generate images grounded in text descriptions. Lastly, we discuss Diffusion Transformer [32], a model that replaces the U-net architecture with a transformer and achieves higher-quality image generation.
@@
$q(x_{t}\mid x_{t-1})$ $=\mathcal{N}\left(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}\mathbf{I}\right)\,,$
$p_{\theta}(x_{t-1}\mid x_{t})$ $=\mathcal{N}\left(x_{t-1};\mu_{\theta}(x_{t},t),\Sigma_{\theta}(x_{t},t)\right)\,,$
@@
Diffusion Transformer
Diffusion Transformer leverages the Vision Transformer [5] structure to replace the U-net architecture traditionally used in diffusion models for noise prediction. Its training and sampling methods remain consistent with DDIM, with the only difference being the replacement of noise prediction network $\epsilon_{\theta}$ with $\epsilon_{\tilde{\theta}}$, where $\tilde{\theta}$ represents a Vision Transformer-based architecture. This approach further enhances the generation quality and ensured that the model possesses good scalability properties.
@@ 4 Algorithm Design
In this section, we introduce our algorithm. We begin by discussing the definition of variation API and the limitations of previous membership inference attack methods. In our formulations, we assume DDIM as our target model. The formulations for DDPM are highly similar and we omit it for brevity. We will discuss the generalization to the latent diffusion model in Section 4.3.
4.1 The variation API for Diffusion Models
Most previous works on membership inference attacks against diffusion models aim to prevent data leakage and hence rely on thresholding the model’s training loss. For instance, [15] involves a direct comparison of image losses, while [6, 19] evaluates the accuracy of the model’s noise prediction at initial or intermediate steps. However, the required access to the model’s internal U-net structure prevents applications from copyright protection because most servers typically provide only black-box API access.
In contrast, our method represents a step towards black-box MIA, as we do not directly access the model’s internal structure. Instead, we rely solely on the variation API, which takes an input image and returns the corresponding output image. Below, we formalize the definition of the variation API used in our algorithm.
Definition 1 (The variation API).
@@ 4.1 The variation API for Diffusion Models
This definition aligns with the image-to-image generation method of the diffusion model, making access to this API practical in many setups [26, 36, 48]. Some APIs provide the user with a choice of $t$, while others do not and use the default parameter. We will discuss the influence of different diffusion steps in Section 5.5, showing that the attack performances are relatively stable and not sensitive to the selection of $t$. We also note that for the target model, we can substitute $\phi_{\theta}(x_{t},0)$ with other sampling methods, such as the Euler-Maruyama Method [28] or Variational Diffusion Models [17].
Figure 1: The overview of ReDiffuse. We independently input the image to the variation API $n$ times with diffusion step $t$. We take the average of the output images and compare them with the original ones. If the difference is below a certain threshold, we determine that the image is in the training set.
@@ 4.2 Algorithm
[…]
$L(\theta)=\mathbb{E}_{\epsilon\sim\mathcal{N}(0,\mathbf{I})}\left[\left\|\epsilon-\epsilon_{\theta}\left(\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,t\right)\right\|^{2}\right]\,.$
Denote $x_{t}=\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon$, we assume that the denoise model is expressive enough such that for the input $x_{0}\in\mathbb{R}^{d}$ and time step $t\in\mathcal{T}$, the Jacobian matrix $\nabla_{\theta}\epsilon_{\theta}(x_{t},t)$ is full rank $(\geq d)$. This suggest that the model can adjust the predicted noise $\epsilon_{\theta}(x_{t},t)$ locally in any direction. Then for a well trained model, we would have
$\nabla_{\theta}L(\theta)=0$ $\implies\nabla_{\theta}\epsilon_{\theta}(x_{t},t)^{T}\mathbb{E}_{\epsilon\sim\mathcal{N}(0,\mathbf{I})}\left[\epsilon-\epsilon_{\theta}\left(\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,t\right)\right]=0\,,$
$\implies\mathbb{E}_{\epsilon\sim\mathcal{N}(0,\mathbf{I})}\left[\epsilon-\epsilon_{\theta}\left(\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,t\right)\right]=0\,.$
Intuitively, this is because if the noise prediction from the neural network exhibited high bias, the network could adjust to fit the bias term, further reducing the training loss.
Figure 2: The intuition of our algorithm design. We denote $x$ as the target image, $\hat{x}_{i}$ as the $i$-th image generated by the variation API, and $\hat{x}$ as the average image of them. For member image $x$, the difference $v=x-\hat{x}$ will be smaller after averaging due to $x_{i}$ being an unbiased estimator.
Therefore, for images in the training set, we expect the network to provide an unbiased noise prediction. Since the noise prediction is typically inaccessible in practical applications, we use the reconstructed sample $\hat{x}$ as a proxy. Leveraging the unbiasedness of noise prediction, we demonstrate that averaging over multiple independent reconstructed samples $\hat{x}_{i}$ significantly reduces estimation error (see Theorem 1). On the other hand, for images that are not in the training set, the neural network may not provide an unbiased prediction at these points. We illustrate the intuition in Figure 2.
With the above intuition, we introduce the details of our algorithm. We independently apply the variation API $n$ times with our target image $x$ as input, average the output images, and then compare the average result $\hat{x}$ with the original image. We will discuss the impact of the averaging number $n$ in Section 5.5. We then evaluate the difference between the images using an indicator function:
$f(x)=\mathbf{1}\left[D(x,\hat{x})<\tau\right]\,.$
@@
Algorithm 1 MIA through ReDiffuse
@@
Analysis
We give a short analysis to justify why averaging over $n$ samples in ReDiffuse can reduce the prediction error for training data. We have the following theorem showing that if we use the variation API to input a member $x\sim D_{\text{training}}$, then the error $\|\hat{x}-x\|$ from our method will be small with high probability.
@@
Suppose the DDIM model can learn a parameter $\theta$ such that, for any $x\sim D_{\text{training}}$ with dimension $d$, the prediction error $\epsilon-\epsilon_{\theta}(\sqrt{\bar{\alpha}_{t}}x+\sqrt{1-\bar{\alpha}_{t}}\epsilon,t)$ is a random variable $X=(X_{1},X_{2},\ldots,X_{d})$ with zero expectation and finite cumulant-generating function for each coordinate [8]. Suppose the sampling interval $k$ is equal to the variation API diffusion step $t$. Let $\hat{x}$ be the average of $n$ output images of $x$ using the variation API. Then we have
$\mathbb{P}(\|\hat{x}-x\|\geq\beta)\leq d\exp\left(-n\min_{i}\Psi^{*}_{X_{i}}\left(\frac{\beta\sqrt{\bar{\alpha}_{t}}}{\sqrt{d(1-\bar{\alpha}_{t})}}\right)\right)\,,$
where $\beta>0$ and $\Psi^{*}_{X_{i}}$ is the Fenchel-Legendre dual of the cumulant-generating function $\Psi_{X_{i}}$.
@@
We note that the unbiased assumption on predicted noise can be strong in practice. For experiments, we expect the mean $\|\mu\|$ of predicted noise for trained data is smaller than the mean $\|\mu^{\prime}\|$ from unseen data. As a result, the empirical best choice of $n$ would be determined by the gap $\|\mu^{\prime}\|-\|\mu\|$.
4.3 MIA on Other Diffusion Models
In this section, we discuss how we can generalize our algorithm to other diffusion models. We note that the variation API for stable diffusion is different from DDIM, as it includes the encoder-decoder process. Again we denote $\hat{x}=V_{\theta}(x,t)$, and the details are as follows:
@@ 4.3 MIA on Other Diffusion Models
For the Diffusion Transformer, we define the variation API as $V_{\tilde{\theta}}(x,t)$, where $\tilde{\theta}$ corresponds to the Vision Transformer architecture instead of the U-net. We repeatedly call the variation API and calculate the difference between the original image and the reconstructed image, as done in DDIM.